# Electricity Demand and Grid Impacts of AI Data Centers: Challenges and Prospects

### v3

@@ VI Conclusion
According to the IEA report [8], global data centers consumed around 415 TWh of electricity in 2024, accounting for about 1.5% of total global electricity consumption. The United States was the largest contributor, accounting for 45% of this demand, followed by China at 25% and Europe at 15%. Given their dominant shares of global data center electricity consumption, the cases in the United States, China, and Europe are discussed in greater detail as below; see [8, 165, 166, 167] for more data center demand information in other countries and regions.
@@
The United States is among the largest markets for hyperscale data centers and accounts for a significant share of global AI-related electricity demand. U.S. data center energy consumption, which reached 176 TWh (4.4% of total U.S. electricity) in 2023, is projected to increase to a range of 325-580 TWh by 2028, contingent upon various growth scenarios [168]. McKinsey & Company projects that U.S. data center load will rise from 25 GW in 2024 to over 80 GW by 2030, largely driven by AI workloads [24]. Moreover, AI data center deployment in the U.S. is highly concentrated, as evidenced by the fact that roughly 80% of the power demand is located in only 15 states [14], with Virginia, Texas, Georgia, and Arizona being leading hubs due to their favorable tax incentives, existing fiber networks, and relatively lower energy costs [169, 170]. Northern Virginia’s “Data Center Alley” alone accounts for more than 2.5 GW of active capacity, with multiple AI-focused data center campuses under construction. New projects are increasingly targeting Midwestern and Mountain West states such as Iowa, Ohio, and Wyoming, for proximity to renewable energy resources and to diversify geographic risk.
@@
China is rapidly expanding its AI data center infrastructure, leading to a commensurate increase in power demand. In 2024, the electricity consumption of data centers in China is estimated to be in the 100-150 TWh range, representing about 1-2% of the nation’s total electricity use, which is projected to rise to between 400 and 600 TWh annually by 2030 [171, 172]. To manage this growth, China has integrated AI development into its national energy strategy. A key strategy is the Eastern Data, Western Computing initiative, which encourages siting new high-performance computing facilities in western provinces like Qinghai, Xinjiang, and Heilongjiang that are rich in renewable resources. This policy, coupled with accelerated smart grid integration for demand forecasting and load management, aims to align the growing computational demand with sustainable power generation and grid reliability [171].
@@
Europe is also experiencing a significant expansion of its AI data center capacity. The total IT load demand for data centers in Europe is projected to increase from approximately 10 GW in 2024 to 35 GW by 2030, with corresponding annual power consumption nearly tripling from 62 TWh to over 150 TWh by 2030, which would represent 5% of the continent’s total electricity use [173]. This growth is underpinned by strategic EU initiatives, including the EUR 30 billion gigafactory plan and the Climate Neutral Data Center Pact, which promote the development of large-scale facilities powered by renewable energy [174]. Projects such as OpenAI’s proposed Stargate facility in Norway, planned to run on hydropower with an initial 230 MW capacity expandable to 520 MW, exemplify this trend [175]. Accommodating this expansion will require an estimated EUR 250-300 billion in grid upgrades and low-carbon energy integration [176]. This investment is necessary to support McKinsey’s forecast of nearly tripled data center electricity consumption by 2030 [173].

### v4

@@ VI Conclusion
According to the IEA report [iea2024energyai], global data centers consumed around 415 TWh of electricity in 2024, accounting for about 1.5% of total global electricity consumption. The United States was the largest contributor, accounting for 45% of this demand, followed by China at 25% and Europe at 15%. Given their dominant shares of global data center electricity consumption, the cases in the United States, China, and Europe are discussed in greater detail as below; see [iea2024energyai, kamiya2025data, ember2025grids, icis2024hungerpwr] for more data center demand information in other countries and regions.
@@
The United States is among the largest markets for hyperscale data centers and accounts for a significant share of global AI-related electricity demand. U.S. data center energy consumption, which reached 176 TWh (4.4% of total U.S. electricity) in 2023, is projected to increase to a range of 325-580 TWh by 2028, contingent upon various growth scenarios [shehabi20242024]. McKinsey & Company projects that U.S. data center load will rise from 25 GW in 2024 to over 80 GW by 2030, largely driven by AI workloads [mckinsey2024datacenters]. Moreover, AI data center deployment in the U.S. is highly concentrated, as evidenced by the fact that roughly 80% of the power demand is located in only 15 states [aljbour2024powering], with Virginia, Texas, Georgia, and Arizona being leading hubs due to their favorable tax incentives, existing fiber networks, and relatively lower energy costs [reuters2025grid, enconnex2024power]. Northern Virginia’s “Data Center Alley” alone accounts for more than 2.5 GW of active capacity, with multiple AI-focused data center campuses under construction. New projects are increasingly targeting Midwestern and Mountain West states such as Iowa, Ohio, and Wyoming, for proximity to renewable energy resources and to diversify geographic risk.
@@
China is rapidly expanding its AI data center infrastructure, leading to a commensurate increase in power demand. In 2024, the electricity consumption of data centers in China is estimated to be in the 100-150 TWh range, representing about 1-2% of the nation’s total electricity use, which is projected to rise to between 400 and 600 TWh annually by 2030 [scmp2024datacenters, carbonbrief2024china]. To manage this growth, China has integrated AI development into its national energy strategy. A key strategy is the Eastern Data, Western Computing initiative, which encourages siting new high-performance computing facilities in western provinces like Qinghai, Xinjiang, and Heilongjiang that are rich in renewable resources. This policy, coupled with accelerated smart grid integration for demand forecasting and load management, aims to align the growing computational demand with sustainable power generation and grid reliability [scmp2024datacenters].
@@
Europe is also experiencing a significant expansion of its AI data center capacity. The total IT load demand for data centers in Europe is projected to increase from approximately 10 GW in 2024 to 35 GW by 2030, with corresponding annual power consumption nearly tripling from 62 TWh to over 150 TWh by 2030, which would represent 5% of the continent’s total electricity use [reuters2024europe]. This growth is underpinned by strategic EU initiatives, including the EUR 30 billion gigafactory plan and the Climate Neutral Data Center Pact, which promote the development of large-scale facilities powered by renewable energy [climatepact2024]. Projects such as OpenAI’s proposed Stargate facility in Norway, planned to run on hydropower with an initial 230 MW capacity expandable to 520 MW, exemplify this trend [itpro2024stargate]. Accommodating this expansion will require an estimated EUR 250-300 billion in grid upgrades and low-carbon energy integration [Cremona2025Grids]. This investment is necessary to support McKinsey’s forecast of nearly tripled data center electricity consumption by 2030 [reuters2024europe].